OpenStudy (anonymous):

Solve over the interval [0, 360): 2sin^2(x)=3-4cos(x) Currently stuck up to: 2cos^2(x)-4cos(x)+1=0

OpenStudy (anonymous):

hint: sin^2x = 1 - cos^2x

OpenStudy (anonymous):

Right, I was able to approach the problem using the Pyth Ident until I got stuck at:

OpenStudy (anonymous):

2cos^2(x)-4cos(x)+1=0

OpenStudy (tkhunny):

Can you solve this? \(2z^{2} - 4z + 1 = 0\)